OpenStudy (hellokitty17):

If -nx2 + tx + c = 0, what is x equal to? x is equal to minus t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n x is equal to t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n x is equal to t plus or minus square root of t squared minus 4 multiplied by n multiplied by c, all divided by minus 2 times n x is equal to minus t plus or minus square root of t squared plus 4 multiplied by n multiplied by c, all divided by minus 2 times n

OpenStudy (hellokitty17):

@MeganChase @jcoury

oregonduck (oregonduck):

Since you have unknown coefficients of x, you're only able to express it in terms of its coefficients (or, as shown below, as 0). Therefore: x = (-t ± √(t²-4n*0))/-2n = (-t±t)/-2n so x = (-t+t)/-2n = 0 or x = (-t-t)/-2n = -2t/-2n = t/n

OpenStudy (anonymous):

Since you have unknown coefficients of x, you're only able to express it in terms of its coefficients (or, as shown below, as 0). Therefore: x = (-t ± √(t²-4n*0))/-2n = (-t±t)/-2n so x = (-t+t)/-2n = 0 or x = (-t-t)/-2n = -2t/-2n = t/n

OpenStudy (anonymous):

wow

oregonduck (oregonduck):

XD LOL

OpenStudy (anonymous):

Sources: My good friends at brainly.com XD

OpenStudy (anonymous):

http://openstudy.com/study#/updates/55ccd0b4e4b0554d62718231 lol noones answering my question

Nnesha (nnesha):

LOLOLOL

Nnesha (nnesha):

if you want to copy the solution from the google so whats the point of asking question here ? :P js

OpenStudy (anonymous):

i didn't ask the question @Nnesha

Nnesha (nnesha):

not you

oregonduck (oregonduck):

that is the way @HelloKitty17 asked me to do it

oregonduck (oregonduck):

medal me ;)